# Beyond the Request: Harnessing HTTP Response Headers for Cross-Browser Web Tracker Classification in an Imbalanced Setting

### v1

@@ Abstract.
The World Wide Web’s connectivity is greatly attributed to the HTTP protocol, with HTTP messages offering informative header fields that appeal to disciplines like web security and privacy, especially concerning web tracking. Despite existing research employing HTTP/S request messages to identify web trackers, HTTP/S response headers are often overlooked. This study endeavors to design effective machine learning classifiers for web tracker detection using HTTP/S response headers. Data from the Chrome, Firefox, and Brave browsers, obtained through the traffic monitoring browser extension T.EX, serves as our data set. Eleven supervised models were trained on Chrome data and tested across all browsers. The results demonstrated high accuracy, F1-score, precision, recall, and minimal log-loss error for Chrome and Firefox, but subpar performance on Brave, potentially due to its distinct data distribution and feature set. The research suggests that these classifiers are viable for detecting web trackers in Chrome and Firefox. However, real-world application testing remains pending, and the distinction between tracker types and broader label sources could be explored in future studies.
Web Privacy, Web Tracking, Web Privacy Measurement, HTTP/S Response Header, Machine Learning
@@ 1. Introduction
The rapid evolution of the Web has cemented its role as an integral component of modern life. Through desktop and mobile browsers, users access a vast array of information and services. However, every interaction, from content consumption to mere site visitation, reveals information about the habits and preferences of users, which in turn has implications for personal privacy and online security.
These privacy concerns are highlighted by the ubiquity of behavioural targeting. By analyzing a user’s browsing history, a detailed profile can be constructed, facilitating the delivery of tailored advertisements and content. This profiling is made possible by web tracking that employs sophisticated techniques to track users across the Web. Given the increasing concerns about such practices (Melicher et al., 2016; Chanchary and Chiasson, 2015), there is a growing emphasis on research in web privacy.
An established countermeasure against unsolicited HTTP communications is the use of filter lists (Traverso et al., 2017), such as EasyList (eas, 2023a) and EasyPrivacy (eas, 2023b). However, their dependency on manual curation makes them vulnerable to oversights and rapid obsolescence (Hashmi et al., 2019; Alrizah et al., 2019; Fouad et al., 2018). Moreover, some trackers employ strategies like domain masking to evade these lists (Chen et al., 2021).
In an attempt to mitigate these shortcomings, recent efforts have explored the potential of Machine Learning (ML) in this domain. Various classifiers, some achieving accuracy metrics in the 80th percentile (Cozza et al., 2020; Iqbal et al., 2021), have been developed using diverse data sources such as HTTP headers and JavaScript calls. Yet, a majority of these focused primarily on HTTP requests, potentially neglecting other informative data streams. Additionally, the predominant reliance on Firefox for data collection may limit the comprehensiveness of these studies.
Notably, there seems to be a gap in literature regarding the use of HTTP response headers as the sole data source for tracker classification across diverse browsers, despite studies in the related field of online security highlighting the efficacy of this approach (McGahagan et al., 2019). This study intends to bridge this gap. We strive to develop classifiers adept at web tracker detection, while confronting challenges such as data sparsity, high dimensionality, and class imbalance.
Our key contributions are:
@@
    We conduct a rigorous study of HTTP response headers across multiple browsers, focusing on their potential for web tracker detection within the Tranco top 10k websites. Furthermore, we evaluate dataset similarities to discern differences in HTTP response headers across the assessed browsers.
@@
    We present a semi-automated ML pipeline which facilitates the training of eleven classifiers tailed for web tracker detection. Our findings indicate that selected tree and gradient boosting models, when applied to a subset of binary-encoded HTTP/S headers, are capable of achieving ROC-AUC, AUPCR, and F1 scores exceeding $0.9$.
@@
    We conduct a comprehensive and multi-faceted evaluation of these classifiers, which includes longitudinal studies to gauge potential concept-drift and evaluates the cross-browser performance of classifiers trained solely on traffic data collected within the Chrome browser. Notably, our evaluation incorporates thirteen metrics, offering a holistic view of the binary classification challenge.
@@
    We assess calibration effects on final classifier performance — a consideration seldom addressed in prior research. This allows us to discern if models trained solely on data from one browser can be effectively deployed on data from alternate browsers without the necessity for retraining.
The remainder of this paper is structured as follows. Section 2 reviews past research in the field of web tracker detection. Section 3 introduced definitions, our research questions, and the analyzed data sets. Section 4 explores HTTP/S response headers from Chrome, Firefox, and Brave data sets, providing general findings and identifies potential challenges for the ML pipeline. Section 5 describes the design and implementation of the ML pipeline, including selected models, metrics, and baselines. Section 6 presents the evaluation of the trained classifiers. The penultimate Section 7 discusses our results and limitations. Section 8 concludes the paper.
@@ 2.1. Classification Based on HTTP Information
Approaches to the detection web trackers with the support of machine learning date back to the early 2000s. In 2010, Yamada et al. (Yamada et al., 2010) were the first to use temporal link analysis to build a graph based on hosts related to HTTP requests from the traffic generated by corporate network. Their classifier considers temporal information such as the visiting time of a specific website or host. Yamada et al. consider this a discriminating criterion based on the assumption that users spend less time on websites of web trackers, as trackers generally appear as third parties that are contacted briefly in the context of other websites rather than first parties that users actively visit.
In 2013, Bau et al. (Bau et al., 2013) described and discussed the usefulness of machine learning as an effective counter to web tracking. They outlined the problems with manually curated filter lists - problems that still hold true today, such as the reliance on manual curation and the inflexibility of static lists. Beyond building an Elastic Net model focused on traces of content extraction in the Document Object Model and HTTP headers to show the advantages of automated approaches, Bau et al. laid down requirements for future machine learning models as well as problems facing researchers in this domain.
@@
In 2015, Gugelmann et al. (Gugelmann et al., 2015) proposed an alternative approach by deriving several statistical features from HTTP traffic, e.g. up- and download volume. In contrast to prior studies, they thoroughly described their model-building process and emphasized their feature engineering. Their feature selection process included the use of Pearson’s correlation coefficient as well as the consideration of a feature’s entropy. In addition, they employed an unspecified classification tree, among other classifiers, to find an optimal Naïve Bayes model that achieved a precision of 83% and was able to correctly identify previously unseen trackers.
@@
In the same year, Li et al. (Li et al., 2015) proposed TrackAdvisor, a system to classify HTTP requests based on the collective statistics of cookies contained therein. Using FourthParty, they captured 563,031 requests and identified 22,270 cookies that were sent to third parties. Rather than relying solely on filter lists, they used a multi-step process with an emphasis on manual inspection to label the captured data. Notably, the training and test sets created from the captured data comprised only 500 requests each and had a balanced class distribution, which contrasts with the distribution found in the original data. Nevertheless, Li et al. used these sets to train a SVM that ultimately achieved a precision of 99.4% and a recall of 100%. Like Gugelmann et al., they point out the usefulness of the HTTP Referrer header for third-party tracker detection.
@@
In 2016, Dudykevych and Nechypor (Dudykevych and Nechypor, 2016) also used cookies as the basis for their classifiers. In contrast to Li et al., however, they developed their own crawler, had a larger data set of 550 000 cookies, and developed a different feature set. The latter is only described briefly, but information from the cookies themselves and the HTTP requests were used as features. It can be assumed that the features are numerical and categorical, as the authors used Principal Component Analysis and correlation, which require numerical data, as well as binary encoding of categorical features. The resulting classification accuracies ranged from 82.01% for Naïve Bayes to 95.43% for the regularized logistic regression method.
@@
The analysis of HTTP traffic for and with machine learning models is also an important topic in web security research, albeit with a different target (i.e. label) than web privacy research. In a 2019 study, McGahagan et al. (McGahagan et al., 2019) analyzed response headers from 45 874 websites and extracted 672 features, resulting in a final feature vector with 22 properties. In addition, they tested whether over- and undersampling had any effects on the model’s performance. The coverage of metrics, models, different training and tuning combinations, and extensive feature engineering documentation can be considered standard among machine-learning practitioners but is somewhat rare among prior related research. An important contribution of this paper is the in-depth presentation of the usefulness of HTTP response headers for automated ML models.
@@
Cozza et al. (Cozza et al., 2020) developed a system that not only classifies HTTP requests but also automatically updates the filter list. Although this study is not the first to detect previously undocumented trackers using machine learning techniques, the deployment of a tracker-detecting browser extension that combines a filter list with a machine learning models and updates its ground truth marks an important step forward in tracker detection technology.
Guarino et al. (Guarino. et al., 2020) built upon prior research by Gugelmann et al. and Li et al. by training multiple model based on combinations of features used by these studies, including different HTTP headers and the ratio of upstream and downstream data transmissions. Their top performer is a Random Forest model with an average accuracy of 92%. It should be noted, however, that this study is based on a relatively small dataset of just 1000 observations. Nevertheless, the authors astutely emphasize that content delivery networks (CDNs) are a primary cause of false positives in the domain of web tracker detection due to similarities in traffic patterns between CDNs and third-party trackers.
Taking advantage of the observation that HTTP-based communication between clients and servers often comprises cascades of multiple HTTP messages, Iqbal et al. (Iqbal et al., 2022) conducted an extensive study that uses entire request chains rather than individual requests as the basis for classifying web tracking. Notably, they were among the first to consider responses as well as requests by developing three distinct feature sets: the first is based on the entire request chain, the second relies on information pertaining to individual requests and the third considers response headers and metadata. By training a Random Forest model on five distinct datasets with different configurations from separate crawls, Iqbal et al. achieved accuracies above 98%. However, if the model was trained on only one crawl and tested on the others, the accuracy dropped by up to 14%, highlighting the difficulties of applying models to datasets with varying configurations, e.g. data collected across different browsers.
@@ 2.2. Other Detection Approaches
Beyond the HTTP-focused approaches outlined above, a variety of classification methods utilizing different data sources for the extraction of distinctive criteria have been proposed and evaluated. These include, classifiers focused on JavaScript code and API access events that aim to identify device fingerprinting techniques, pure URL-based classifiers, CNAME cloaking detectors, and graph-based classifiers that leverage the topology of the Web to identify third-party tracking in particular. In the following paragraphs, we will briefly highlight notable research on these alternative approaches.
Focusing on JavaScript API calls, Wu et al. (Wu et al., 2016) developed DMTTrackerDetector, a system using decision trees in which observed API calls serve as features. The system, comprising a 717-dimensionial feature vector, achieved 97.8% accuracy and addressed obfuscation issues, showing the distinct JavaScript API usage by trackers and non-trackers.
Bahrami et al. (Bahrami et al., 2021) refined the scope of their analysis by focusing on device fingerprinting rather than API calls in general. Their approach combines graph analysis with supervised and unsupervised learning to detect web trackers in historical JavaScript file data collected between 2010 and 2019 by creating a graph that includes temporal aspects and the output of AST-based keyword extraction. Their method, which includes nine graph-related features and graph embedding, as well as a clustering model, yielded accuracies up to 89.13% and identified distinct clusters for fingerprinting and functional APIs, highlighting the viability of such an approach.
@@
Combining these graph-based approaches, Castell-Uroz et al. (Castell-Uroz et al., 2021) proposed a tripartite network graph to identify third-party tracking resources and their origin domains within first-party websites. By applying this approach to a dataset comprising over 90 thousand domains, they were able to detected new trackers through their hash value popularity and dirt level, finding that higher values correlated with the likelihood of a given resource to be tracking-related.
@@
Finally, Yu et al. (Yu et al., 2016) proposed a browser extension using k-Anonymity principles to classify third-party trackers based on data sent to users. Leveraging data from 200,000 German users of the Cliqz browser, their system outperformed traditional blockers like Disconnect, which had a 38.8% false positive rate.
@@ 2.3. Summary
Despite over a decade of dedicated research, web tracking is as prevalent as ever, new tracking methods are developed, and the most widespread form of protection, namely blockers and their filter lists, exhibit several key problems. Filter lists, which are also used as the ground truth for labeling in many studies related to web privacy, are often maintained by a small group of users, contain irrelevant rules, are rather slow to add new rules, and can simply be circumvented by changing URLs. In response, researchers have developed a wealth of automated approaches, utilizing HTTP messages, JavaScript, or the HTML files of websites themselves to classify third parties as trackers.
A crucial component of this ongoing research effort is the availability and quality of data on which to train and evaluate machine learning approaches.
Despite this, HTTP response headers are often neglected in this regard. There are cases where authors document the collection of HTTP responses (Acar et al., 2013) but not their actual use. In other cases, the data is collected, but only used in support of other approaches such as the classification of cookies (Lerner et al., 2016; Englehardt et al., 2015).
Additionally, the majority of studies are based on datasets collected using just one browser, most commonly Firefox. Cassel et al. (Cassel et al., 2022) discuss the need for cross-browser examinations to avoid biased datasets. This is supported by research that shows that characteristics such as JavaScript API calls differ between browsers (Yang and Yue, 2020).
In summary, our review of related research has shown that, although the application of machine learning to the field of web tracking identification and classification has been investigated in numerous studies, a research gap still exists with regard to cross-browser performance of classifiers and the usefulness of HTTP response data for classification purposes.

### v2

@@ Abstract.
The World Wide Web’s connectivity is greatly attributed to the HTTP protocol, with HTTP messages offering informative header fields that appeal to disciplines like web security and privacy, especially concerning web tracking. Despite existing research employing HTTP request messages to identify web trackers, HTTP response headers are often overlooked. This study endeavors to design effective machine learning classifiers for web tracker detection using binarized HTTP response headers. Data from the Chrome, Firefox, and Brave browsers, obtained through the traffic monitoring browser extension T.EX, serves as our dataset. Ten supervised models were trained on Chrome data and tested across all browsers, including a Chrome dataset from a year later. The results demonstrated high accuracy, F1-score, precision, recall, and minimal log-loss error for Chrome and Firefox, but subpar performance on Brave, potentially due to its distinct data distribution and feature set. The research suggests that these classifiers are viable for web tracker detection. However, real-world application testing remains pending, and the distinction between tracker types and broader label sources could be explored in future studies.
Web Privacy, Web Tracking, Web Privacy Measurement, HTTP Response Header, Machine Learning
@@ 1. Introduction
The rapid evolution of the Web has cemented its role as an integral component of modern life. Users access a vast array of information and services through desktop and mobile browsers. However, every interaction, from content consumption to mere site visitation, reveals information about the habits and preferences of users, which in turn has implications for personal privacy and online security.
These privacy concerns are highlighted by the ubiquity of behavioural targeting. A detailed behavioural profile can be constructed by analyzing a user’s browsing history, facilitating the delivery of tailored advertisements and content. This profiling is made possible by web tracking that employs sophisticated techniques to track users across the Web. Given the increasing concerns about such practices (Melicher et al., 2016; Chanchary and Chiasson, 2015), there is a growing emphasis on research in web privacy.
An established countermeasure against unsolicited HTTP communications is the use of filter lists (Traverso et al., 2017) such as EasyList  [^1] https://easylist.to/easylist/easylist.txt (EL) and EasyPrivacy  [^2] https://easylist.to/easylist/easyprivacy.txt (EP). However, their dependency on manual curation makes them vulnerable to oversights and rapid obsolescence (Hashmi et al., 2019; Alrizah et al., 2019; Fouad et al., 2018). Moreover, some trackers employ strategies like domain masking to evade these lists (Chen et al., 2021).
In an attempt to mitigate these shortcomings, efforts of past decades have explored the potential of Machine Learning (ML) in this domain. Various classifiers, some achieving accuracy metrics in the 80th percentile (Cozza et al., 2020; Iqbal et al., 2021), have been developed using diverse data sources such as HTTP headers and JavaScript API calls.
Meanwhile, the potential of information contained in HTTP responses has been largely overlooked, except for HTTP cookies. The value of this information might be doubted because it becomes available only after web tracking has occurred. However, this limitation primarily affects scenarios where the model’s predictions are used to block communications with web trackers in real-time. Alternative approaches are feasible: (i) dynamically generating filter lists based on the model’s predictions, or (ii) using it as a supplementary tool for classifiers that depend solely on pre-HTTP request information.
Most previously proposed classifiers have not addressed the issue of deployability, suggesting that the field is still in an exploratory phase. Consequently, it is valuable to examine the suitability of HTTP response information for web tracker classification. Unlike HTTP request information, which is controlled by the client and varies between users, HTTP response information is configured by the server, i.e., the potential web tracker. Notably, a study in the related field of web security underscores the effectiveness of this approach (McGahagan et al., 2019).
Our key contributions are as follows:
@@
    We conduct a study of HTTP response headers across multiple browsers, focusing on their potential for web tracker detection within the Tranco global top 10K websites. Furthermore, we identify challenges with this type of data for developing ML classifiers.
@@
    We present a semi-automated ML pipeline which facilitates the training of ML models for the binary classification of web trackers based on HTTP response or request headers.
@@
    We train and evaluate ten supervised classifiers. Our findings indicate that when applied to a subset of binarized HTTP headers, selected tree and gradient boosting models can detect web trackers. Our best classifier achieves ROC-AUC, AUPRC, and an F1-score exceeding $0.93$.
@@
    We conduct a comprehensive and multi-faceted evaluation of these classifiers, which includes longitudinal studies to gauge potential concept drift and evaluates the cross-browser performance of classifiers trained solely on traffic data collected within the Chrome browser. We show that Chrome data captures a majority of the structural header characteristics of trackers. Our evaluation incorporates 13 metrics, offering a holistic view of the binary classification challenge. In addition, we use our pipeline to train classifiers using only HTTP request headers to highlight the comparative value of HTTP response headers for web tracker detection.
The remainder of this paper is structured as follows: Section 2 reviews past research in the field of web tracker detection. Section 3 introduced definitions, our research questions, and the analyzed datasets. Section 4 explores HTTP response headers from Chrome, Firefox, and Brave datasets, providing general findings and identifying potential challenges for the ML pipeline. Section 5 describes the design and implementation of the ML pipeline, including selected models, metrics, and baselines. Section 6 presents the evaluation of the trained classifiers. The penultimate Section 7 discusses our results and limitations, and Section 8 concludes the paper.
@@ 2.1. Classification Based on HTTP Information
Approaches to the detection of web trackers with the support of machine learning date back to the early 2000s. In 2010, Yamada et al. (Yamada et al., 2010) were the first to use temporal link analysis to build a graph based on hosts related to HTTP requests from the traffic generated by corporate networks. Their classifier considers temporal information, such as the visit duration at a specific website or host. Yamada et al. consider this a discriminating criterion based on the assumption that users spend less time on websites of web trackers, as trackers generally appear as third parties that are contacted briefly in the context of other websites rather than first parties that users actively visit.
In 2013, Bau et al. (Bau et al., 2013) described and discussed the usefulness of machine learning as an effective counter to web tracking. They outlined problems with manually curated filter lists - problems that still hold true today, such as the reliance on manual curation and the inflexibility of static lists. Beyond building an Elastic Net model focused on traces of content extraction in the Document Object Model and HTTP headers to show the advantages of automated approaches, Bau et al. laid down requirements for future machine learning models as well as problems facing researchers in this domain.
@@
In 2015, Gugelmann et al. (Gugelmann et al., 2015) proposed an alternative approach by deriving several statistical features from HTTP traffic, e.g., up- and download volume. In contrast to prior studies, they thoroughly described their model-building process and emphasized their feature engineering. Their feature selection process included the use of Pearson’s correlation coefficient as well as the consideration of a feature’s entropy. In addition, they employed an unspecified classification tree, among other classifiers, to find an optimal Naïve Bayes model that achieved a precision of 83% and was able to correctly identify previously unseen trackers.
@@
In the same year, Li et al. (Li et al., 2015) proposed TrackAdvisor, a system to classify HTTP requests based on the collective statistics of cookies contained therein. Using FourthParty, they captured 563,031 requests and identified 22,270 cookies sent to third parties. Rather than relying solely on filter lists, they used a multi-step process emphasising manual inspection to label the captured data. Notably, the training and test sets created from the captured data comprised just 500 requests each and had a balanced class distribution, which differs from the distribution found in the original data. Nevertheless, Li et al. used these sets to train an SVM that ultimately achieved a precision of 99.4% and a recall of 100%. Like Gugelmann et al., they point out the usefulness of the HTTP Referrer header for third-party tracker detection.
@@
In 2016, Dudykevych and Nechypor (Dudykevych and Nechypor, 2016) also used cookies as the basis for their classifiers. In contrast to Li et al., however, they developed their own crawler, used a larger dataset of 550,000 cookies, and developed a different feature set. The latter is only briefly described, but information from the cookies and HTTP requests were used as features. It can be assumed that the features are numerical and categorical, as the authors used Principal Component Analysis and correlation, which require numerical data, as well as binarization of categorical features. The resulting classification accuracies ranged from 82.01% for Naïve Bayes to 95.43% for the regularized logistic regression method.
@@
The analysis of HTTP traffic for and with machine learning models is also an important topic in web security research, albeit with a different target (i.e., label) than web privacy research. In a 2019 study, McGahagan et al. (McGahagan et al., 2019) analyzed response headers from 45 874 websites and extracted 672 features, resulting in a final feature vector with 22 properties. In addition, they tested whether over- and undersampling affected the model’s performance. Their extensive coverage of metrics, models, different training and tuning combinations, and feature engineering documentation can be considered standard among ML practitioners but is somewhat rare among related research in the field of web tracking detection. Crucially, an important contribution of this paper is the in-depth presentation of the usefulness of HTTP response headers for ML models to detect malicious entities. We apply a similar approach but to web privacy research.
@@
Cozza et al. (Cozza et al., 2020) developed a system that not only classifies HTTP requests but also automatically updates the filter list. Although this study is not the first to detect previously undocumented trackers using machine learning techniques, the deployment of a tracker-detecting browser extension that combines a filter list with a machine learning model and updates its ground truth marks an important step forward in tracker detection technology.
Guarino et al. (Guarino. et al., 2020) built upon prior research by Gugelmann et al. and Li et al. by training multiple models based on combinations of features used by these studies, including different HTTP headers and the ratio of upstream and downstream data transmissions. Their top performer is a Random Forest model with an average accuracy of 92%. It should be noted, however, that this study is based on a relatively small dataset of just 1000 observations. Nevertheless, the authors astutely emphasize that content delivery networks (CDNs) are a primary cause of false positives in the domain of web tracker detection due to similarities in traffic patterns between CDNs and third-party trackers.
Taking advantage of the observation that HTTP-based communication between clients and servers often comprises cascades of multiple HTTP messages, Iqbal et al. (Iqbal et al., 2022) conducted an extensive study that uses entire request chains rather than individual requests as the basis for classifying web tracking. Notably, they were among the first to consider responses alongside requests by developing three distinct feature sets: the first is based on the entire request chain, the second relies on information pertaining to individual requests, and the third considers response headers and metadata. By training a Random Forest model on five distinct datasets with different configurations from separate crawls, Iqbal et al. achieved accuracies above 98%. However, if the model was trained on only one crawl and tested on the others, the accuracy dropped by up to 14%, highlighting the difficulties of applying models to datasets with varying configurations, e.g., data collected across different browsers. Whereas Iqbal et al. selected features based on intuition and domain knowledge, our approach selects features from a data-driven standpoint and considers all response headers for feature extraction.
@@ 2.2. Other Detection Approaches
Beyond the HTTP-focused approaches outlined above, various classification methods utilizing different data sources to extract distinctive criteria have been proposed and evaluated. These include classifiers focused on JavaScript code and API access events that aim to identify device fingerprinting techniques, pure URL-based classifiers, CNAME cloaking detectors, and graph-based classifiers that leverage the topology of the Web to identify third-party tracking in particular. The following paragraphs briefly highlight notable research on these alternative approaches.
Focusing on JavaScript API calls, Wu et al. (Wu et al., 2016) developed DMTTrackerDetector, a system using decision trees in which observed API calls serve as features. Their system, comprising a 717-dimensional feature vector, achieved 97.8% accuracy and addressed obfuscation issues, showing the distinct JavaScript API usage by trackers and non-trackers.
Bahrami et al. (Bahrami et al., 2021) refined the scope of their analysis by focusing on device fingerprinting rather than API calls in general. Their approach combines graph analysis with supervised and unsupervised learning to detect web trackers in historical JavaScript file data collected between 2010 and 2019 by creating a graph that includes temporal aspects and the output of Abstract Syntax Tree-based keyword extraction. Their method, which includes nine graph-related features and graph embedding, as well as a clustering model, yielded accuracies up to 89.13% and identified distinct clusters for fingerprinting and functional APIs, highlighting the viability of such an approach.
@@
Combining these graph-based approaches, Castell-Uroz et al. (Castell-Uroz et al., 2021) proposed a tripartite network graph to identify third-party tracking resources and their origin domains within first-party websites. By applying this approach to a dataset comprising over 90,000 domains, they detected new trackers through their hash value popularity and dirt level, finding that higher values correlated with the likelihood of a given resource to be tracking-related.
@@ 2.3. Summary
Research presented in this section explored various approaches to detect web trackers using machine learning, utilizing various available information. However, HTTP response information has been mostly neglected, and if considered, only the cookie information or a small subset of header information was taken into account. This gap may stem from the intuition that the model’s predictions should be utilized in a real-time application scenario, ignoring use cases where ex-post detection of trackers is sufficient. However, prior studies often owe a thorough discussion on the deployability of their approaches. Related to this, training and test data have been generated by the same browser, often in the same browsing session. The generalization of a model’s predictions is yet to be evaluated. Furthermore, the metrics used to assess a classifier’s performance are not homogenized in the related literature, making a direct comparison of the classifiers’ performances unfeasible.